KynxKatz:

C=5/9(F−32) The equations above shows how temperature F, measured in degrees Fahrenheit, relates to a temperature C, measured in degrees Celsius. Based on the equation, which of the following must be true? I. A temperature increase of 1 degree Fahrenheit is equivalent to a temperature increase of 5/9 degrees Celsius. II. A temperature increase of 1 degree Celsius is equivalent to a temperature increase of 1.8 degrees Fahrenheit. III. A temperature increase of 5/9 degree Fahrenheit is equivalent to a temperature increase of 1 degree Celsius. A. I only. B. II only. C. III only. D. I and II only.

AZ:

Hello

AZ:

Let's take a look at case 1 first We increase the temperature in Fahrenheit by 1, will the temperature in Celsius change by 5/9? let's see C=5/9(F−32) Let's first say that we have 32 Fahrenheit what would you get? C = 5/9 (32- 32) C = ?

AZ:

While I wait and hope that your silence doesn't mean you're confused, @KynxKatz , I shall proceed a little bit further And after you calculate that, let's use 33 Fahrenheit. So that way we can test option 1. What happens when we increase the temperature by 1 F So we have C = 5/9 (33-32) C = ?? Is the difference between the first answer and this one 5/9? If so, that makes condition 1 true

KynxKatz:

Omg I’m so srry

KynxKatz:

I have a lot of questions

AZ:

Sure! We can get through them once we're done with this one :)

KynxKatz:

Mk

KynxKatz:

Nd yes i got the first answer thx

AZ:

You answered the entire question? Or you were able to figure out whether the first I. answer choice was correct?

KynxKatz:

No I answered the whole question

KynxKatz:

I knew it but i just wasnt sure

AZ:

If you want me to check your answer, just let me know!